Charles just finished building an underground pool in his backyard. The rectangular pool has dimensions of 20 feet by 40 feet. He wants to surround it with a pool deck of uniform width. He must use all 16 cubic yards of concrete that were leftover from the pool construction. If one cubic yard of concrete will cover 54 square feet, how wide of a pool deck can he build around the pool?
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

16 cubic yards of concrete that were leftover from the pool construction.
If one cubic yard of concrete will cover 54 square feet,
Area covered = 16*54=864 ft^2
Area of pool deck = total area - area of pool
From fig:
.
(40+2x)(20+2x) -800 = 864
800+80x+40x+4x^2-800 = 864
4x^2+120x-864=0
(x+36)(x-6)=0
ignore negative
x=6
width =6ft

Answer by ikleyn(52878)